OpenStudy (anonymous):

Can Someone please explain the golden rule for solving equations? and if you can please give an example? Thank You!

jimthompson5910 (jim_thompson5910):

the basic idea is to follow PEMDAS in reverse so you can isolate your variable

OpenStudy (anonymous):

i don't know all about them but a few are 1) What you are doing on one side do it on other side eg - x-10 = 5 x-10 + 10 = 5 + 10 [adding 10 on both the sides] x= 15 or 10x = 100 10x/10 = 100/10 x = 10 or x/ 10 = 10 x/10 * 10 = 10 * 10 x = 100 I hope you have understood

jimthompson5910 (jim_thompson5910):

for example say you had 2*3 + 5 = 10 normally you follow PEMDAS to evaluate the left side so it turns into the right side but if you turn the '3' into x, then you get 2x + 5 = 10 to isolate x, you have to undo everything, which means you follow PEMDAS in reverse

OpenStudy (mathstudent55):

This is how I look at it. Look at the variable you are interested in. Notice all the operations being done to it. If you had a number instead of the variable, you would do those operations in a certain order. What you do is, do the opposite operations in the opposite order in order to undo those operations and isolate the variable. Fo example, 5x - 12 = 28 If x were a number you would 1. multiply 5 by x 2. subtract 12 Now do the opposite operation in the opposite order. Start with 2. but in reverse: add 12 Then go on to 1 in reverse: divide by 5 Obviously, what you do to one side you must do to the other side also. So this would be: 5x - 12 = 28 Add 12 to both sides: 5x = 40 Divide both sides by 5 x = 8

OpenStudy (anonymous):

o_o (mind explodes) ty lol

OpenStudy (mathstudent55):

The more you practice, the easier and more second nature it'll become, and then your mind will explode less.